OpenStudy (anonymous):

Express the following function, F(x) as a composition of two functions f and g. F(x) = x^2/(x^2+4) @wio

OpenStudy (anonymous):

Well, the easy way is to say \(g(x) =1\) and \(f(x) = F(x)\).

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Another way is to say \[ F(x) = \frac{x^2}{x^2+4} = \frac{f(x)}{g(x)} \]Which means \[ f(x)=x^2\\ g(x) = x^2+4\\ F =\frac fg \]

OpenStudy (anonymous):

so i just substitute?

OpenStudy (anonymous):

I guess so

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you know how to do it? can you show me?

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

I think I already showed you how to do it.

OpenStudy (anonymous):

Wait, actually

OpenStudy (anonymous):

Okay, I messed up, let me do it again.

OpenStudy (anonymous):

Another way is to say \[ F(x) = \frac{x^2}{x^2+4} = \frac{(x^2)}{(x^2)-4} \]If we let \(g(x) = x^2\): \[ F(x)=\frac{g(x)}{g(x)+4} \]Now, we want \[ f(g) = \frac{g}{g+4} \]Which means that \[ f(x) =\frac{x}{x+4} \]So now we are done because \[ F(x) = (f\circ g)(x) \]

OpenStudy (anonymous):

So now, my final answer is \[x / x + 4\]

OpenStudy (anonymous):

That is just \(f(x)\). Your final answer is providing \(f(x)\) and \(g(x)\).

OpenStudy (anonymous):

ok so how do I get g(x)? We found f(x) already right?

OpenStudy (anonymous):

We already did get \(g(x)\), just look up.

OpenStudy (anonymous):

oh, okay. thanks!

OpenStudy (anonymous):

x/ x + 4 AND g/ g + 4 ARE MY FINAL ANSWERS RIGHT? @wio